OpenStudy (anonymous):

Explain pigeon hole principle. Using this principle show that in any group of 36 people, we can always find 6 people who were born on the same day of week.

OpenStudy (anonymous):

There are 7 days in the week. Pidgeon hole principle says at least one hole (day of the week) that: \[\left\lceil\frac{36}{7} \right\rceil = 6\] will have at least \(6\) pidgeons (birthdays).

OpenStudy (anonymous):

If you have \(p\) pigeons and \(h\) holes, then some hole must have at least \(k\) pidgeons: \[ \left\lceil \frac{p}{h} \right\rceil =k \]